OpenStudy (anonymous):

What is the value of the expression 7 of the power –3? –21 -1/343 1/343 343

OpenStudy (anonymous):

First one since I'm guessing the question was 7 x -3

OpenStudy (anonymous):

oo editted ok =1/7^3 =1/343

HanAkoSolo (jamierox4ev3r):

think of \(7^{3}\) then think of \(7^{-1}\) \(\large 7^{3}=343\) \(\large 7^{-1}=\frac{1}{7}\)

HanAkoSolo (jamierox4ev3r):

a negative exponent switches the numerator and the denominator For example, if \(7^{-1}\) is equal to \(\large\frac{1}{7}\), this means that \(7^{1}\) was equal to \(\large\frac{7}{1}\), which is also 7.

HanAkoSolo (jamierox4ev3r):

so if \(7^{3}\) is equal to \(\large\frac{343}{1}\), then \(7^{-3}\) is equal to? @cakequeen_22